Question 10 of 10, Step 1 of 10 Correct A telephone company claims that the service calls which they receive are equally distributed among the five working days of the week.A survey of 80 randomly selected service calls was conducted. Iis there enough evidence to refute the telephone companys claim that the number of service calls does not change from day-to-day? Days of the Week Mon Tue Wed Thu Fri Number of Calls 19 12 13 17 19 Copy Data Step 1 of 10: State the null and alternative hypothesis Answer Ho: Service calls are not equally distributed over the five working days. Ha: Service calls are equally distributed over the five working days Hn: Service calls are equally distributed over the five working davs o 2018 Hawkes L
Step 1 of 10 : State the null and alternative hypothesis.
Step 2 of 10 : What does the null hypothesis indicate about the proportions of service calls received each day?
Step 3 of 10 : State the null and alternative hypothesis in terms of the expected proportions for each category.
Step 4 of 10 : Find the expected value for the number of service calls received on Monday. Round your answer to two decimal places.
Step 5 of 10 : Find the expected value for the number of service calls received on Wednesday. Round your answer to two decimal places.
Step 6 of 10 : Find the value of the test statistic. Round your answer to three decimal places.
Step 7 of 10 : Find the degrees of freedom associated with the test statistic for this problem.

Step 8 of 10 : Find the critical value of the test at the 0.025
0.025
level of significance. Round your answer to three decimal places.
Step 9 of 10 : Make the decision to reject or fail to reject the null hypothesis at the 0.025
0.025
level of significance.
Step 10 of 10 : State the conclusion of the hypothesis test at the
Answer #1

Step 1 of 10: H0: Null Hypothesis: Service calls are equally distributed over the five working days

                HA: Alternative Hypothesis: Service calls are not equally distributed over the five working days

Step 2 of 10: The null hpothesis indicates the proportion of service calls received each day is same.

Step 3 of 10: H0: Expected proportions for each category is the same

                    HA: Expected proportions for each category is not the same

Step 4 of 10: Expected number of service calls on Monday = 80/5=16.00

Step 5 of 10: Expected number of service calls on Wednesday = 16.00

Step 6 of 10:

O               E                     (O - E)2/E

19              16                      0.5625

12               16                     1

13                16                     0.5625

17               16                     0.0625

19               16                    0.5625

-----------------------------------------------------

\chi ^{2}                           = 2.75

Step 7 of 10:

Degrees of freedom = 5 - 1 = 4

Step 8 of 10:

\alpha = 0.025

From Table, critical value of \chi ^{2} = 11.1433

Step 9 of 10:

Since calculated value of \chi ^{2} = 2.75 is less than critical value of \chi ^{2} = 11.1433, Fail to reject the null hypothesis at the 0.025 level of significance.

Step 10 of 10:

Conclusion:
The data support the claim that the service calls which they reeive are equally distributed among the five working days of the week.